A bicycle leaves town at 16 mph at 8 am. A car leaves at 11 am, traveling in the same direction at 40 mph. How many minutes after 11 am is it when the car overtakes the bicycle?
Answer by ankor@dixie-net.com(15747)

A bicycle leaves town at 16 mph at 8 am. A car leaves at 11 am, traveling in the
same direction at 40 mph. How many minutes after 11 am is it when the car
overtakes the bicycle?
:
Let t = travel time (in hrs) of the car when it overtakes the bike
Then
(t+3) = travel time of the bike
:
We know that when this happens, they will have traveled the same distance.
Make a distance equation from this fact: Distance = Speed * time
:
Car dist = Bike dist
40t = 16(t+3)
40t = 16t + 48
40t - 16t = 48
24t = 48
t = 48/24
t = 2 hrs
:
To answer the question, 120 minutes after 11
:
Check solution by finding the distance of each:
16(2+3) = 80
40(2) = 80 miles also